OpenStudy (anonymous):

HELP ME PLEASE! i will give MEDALS!! its multiple choice! I'm not sure how to solve this. Plans for a bridge are drawn on a coordinate grid. One girder of the bridge lies on the line y= -9x -9. A perpendicular brace passes through the point (-1, 7). Write an equation of the line that contains the brace. The multiple choice is a. y-1= -1/9 (x+7) b. x-7= -9 (y+1) c. y-7= 1/9 (x+1) d. y-7= -9 (x+1) (The 1/9 is a fraction not division!)

OpenStudy (esshotwired):

I believe it is C. Do you want an explanation?

OpenStudy (anonymous):

Yes please thank you so much!

OpenStudy (esshotwired):

For any perpendicular lines, the slopes are always opposite reciprocals. So in this case, the slope of the equation given is -9. You tell me what is the opposite reciprocal of -9?

OpenStudy (anonymous):

First find the slope of y= -9x -9. let us compare it with y=mx+c we find that: slope m1=-9 therefopre slope of the brace passes through the point (-1, 7) would be \[m_2 = \frac{1}{9}\] Now equation of line in slope point form is given as : y-y1=m2(x-x1) y-7=1/9(x-(-1)) y-7=1/9(x+1) is the required equation of the brace. Hence option c. y-7= 1/9 (x+1) is the correct answer. @nialler143

OpenStudy (anonymous):

Thank you both so much! This is very helpful :) I have one more question would one or both of you mind trying to help?

OpenStudy (esshotwired):

Sure. I will see if I can help.

OpenStudy (anonymous):

okay thank you here is it. Write a paragraph proof of this theorem: In a plane, if two lines are perpendicular to the same line, then they are parallel to each other. Given that r is perpendicular to s, t is perpendicular to s.. Prove that r is parallel to t

OpenStudy (esshotwired):

If r is perpendicular to s, then by the definition of perpendicular lines angle 4 has a measure of 90 degrees. Using the same logic with line t, angle 6 is 90 degrees. Since the same side interior angles are supplementary, line r is parallel to line t. I found this image online, and it looked like it went along with this. http://tinypic.com/view.php?pic=rszvdg&s=5#.UlQ43BAixus Hope this helped.

OpenStudy (anonymous):

Yes! thank you again!